Snider Industries sells on terms of 3/10, net 25. Total sales for the year are $1,750,000. Thirty percent of customers pay on the 10th day and take discounts; the other 70% pay, on average, 30 days after their purchases. Assume 365 days in year for your calculations. What is the days sales outstanding? Round your answer to one decimal place. days What is the average amount of receivables? Round your answer to the nearest dollar. Do not round intermediate calculations. $ What would happen to average receivables if Snider toughened its collection policy with the result that all nondiscount customers paid on the 25th day? Round your answer to the nearest dollar. Do not round intermediate calculations. $

Cost of Trade Credit

A large retailer obtains merchandise under the credit terms of 1/15, net 45, but routinely takes 70 days to pay its bills. (Because the retailer is an important customer, suppliers allow the firm to stretch its credit terms.) What is the retailer's effective cost of trade credit? Assume 365 days in year for your calculations. Do not round intermediate calculations. Round your answer to two decimal places.

%
Answer #1

1-a). Days Sales Outstanding = [0.30 * 10] + [0.70 * 30] = 3 + 21 = 24 days

1-b). Average Receivables = DSO * [Sales / 365]

= 24 * [$1,750,000 / 365] = 24 * $4,794.52 = $115,068.49, or $115,069

1-c). Days Sales Outstanding = [0.30 * 10] + [0.70 * 25] = 3 + 17.5 = 20.5 days

Average Receivables = DSO * [Sales / 365]

= 20.5 * [$1,750,000 / 365] = 24 * $4,794.52 = $98,287.67, or $98,288

2). Effective Cost of Trade Credit = [1 + {Discount % / (1 - Discount %)}][365 / (Days Credit Outstanding - Discount Period)] - 1

= [1 + {0.01 / (1 - 0.01)}][365 / (70 -15)] - 1

= [1 + 0.0101][365 / 55] - 1

= [1.0101]6.6364 - 1 = 1.0690 - 1 = 0.0690, or 6.90%